On March 1, Pimlico Corporation (a U.S.-based company) expects to order merchandise from a supplier in Sweden in three months. On March 1, when the spot rate is $0.43 per Swedish krona, Pimlico enters into a forward contract to purchase 567,500 Swedish kroner at a three-month forward rate of $0.440. At the end of three months, when the spot rate is $0.436 per Swedish krona, Pimlico orders and receives the merchandise, paying 567,500 kroner. What amount does Pimlico report in net income as a result of this cash flow hedge of a forecasted transaction?

A. $5,675 Discount expense plus a $1,135 positive adjustment to net income when the merchandise is purchased.

B. $2,270 premium expense plus a $1,135 negative adjustment to net income when the merchandise is purchased.

C. $2,270 premium expense plus a $2,270 positive adjustment to net income when the merchandise is purchased.

D. $5,675 premium expense plus a $3,405 positive adjustment to net income when the merchandise is purchased.
Answer #1

D. $5,675 premium expense plus a $3,405 positive adjustment to net income when the merchandise is purchased.

Premium Expense = 567500(0.44 - 0.43) = 5,675

Positive adjustment to net income = 567500(0.436 - 0.43) = 3,405